# Shell quenching in nuclear charge radii based on Monte Carlo dropout Bayesian neural network

### v1

@@ Abstract
Charge radii can be generally used to encode the information about various fine structures of finite nuclei. In this work, a constructed Bayesian neural network based on Monte Carlo dropout approach is proposed to accurately describe the charge radii of nuclei with proton number $Z\geq 20$ and mass number $A\geq 40$. More motivated underlying mechanisms are incorporated into this combined model except for the basic building blocks with the specific proton and neutron numbers, which naturally contains the pairing effect, isospin effect, shell closure effect associated to the Casten factor $P$, valence neutrons, valence protons, quadrupole deformation $\beta_{20}$, high order hexadecapole deformation $\beta_{40}$, and the local “abnormal” shape staggering effect of 181,183,185Hg. To avoid the over-fitting puzzle along $Z=28$, $50$, and $82$ isotopic chains, the modified Casten factor $P^{*}$ is tentatively introduced into the input structure parameter sets. We have successfully demonstrated the ability of the Monte Carlo dropout Bayesian neural network (MC-dropout BNN) model to significantly increase the accuracy in predicting the nuclear charge radii. The standard root-mean-square deviation falls into $0.0084$ fm for the training data and $0.0124$ fm for the validation data with the modified Casten factor $P^{*}$ input. Meanwhile, the shell closure effect of nuclear charge radii can be reproduced remarkably well.
@@ I Introduction
Finite nuclei, which are characterized as the strongly correlated many-body systems, provide a natural probe to recognize the fundamental interactions and forces. As directly measured quantities in atomic nuclei, charge radii can be generally used to reflect the proton density distributions and encode the information about various fine structure phenomena, such as the emergence of shell closure effects Angeli_2015 ; PhysRevC.100.044310 ; GarciaRuiz:2019cog ; PhysRevC.104.064313 ; PhysRevLett.129.142502 and the odd-even staggering (OES) behavior GarciaRuiz:2016ohj ; Marsh:2018wxs ; Miller2019 ; PhysRevLett.127.192501 . Furthermore, available description of proton density distributions can also be used to pin down the components of isospin interactions in the equation of state (EoS) of asymmetric nuclear matter PhysRevC.88.011301 ; PhysRevLett.119.122502 ; PhysRevResearch.2.022035 ; PhysRevLett.127.182503 ; nuclscitech34.119 ; PhysRevLett.132.162502 . This means that more data about nuclear charge radii should be urgently required with the highly accurate precision. Remarkable achievements have been obtained in detecting charge radii over the past decades. Especially for nuclei with extreme neutron-to-proton ratios, laser spectroscopy method can be considered as an available access to measure the nuclear size CAMPBELL2016127 ; YANG2023104005 .
@@
The undertaken efforts have been devoted to featuring the normal and abnormal variations in nuclear charge radii along a long isotopic family, which includes the empirical formulas Zhang2002EPJA285 ; PhysRevC.87.024310 ; Piekarewicz2010 ; PhysRevC.94.064315 , relativistic mean field theory Geng:2003pk ; PhysRevC.102.024307 ; PhysRevC.109.064302 , non-relativistic Skyrme Hartree-Fock-Bogoliubov (HFB) approach PhysRevLett.102.242501 ; PhysRevC.82.035804 , and the Fayans energy density functional PhysRevC.95.064328 ; PhysRevLett.122.192502 , etc. Recently, machine learning methods have been made greatly applications in the course of nuclear physics  PhysRevC.98.034318 ; Bedaque2021 ; Paquet_2024 . Particularly, a well-known approach based on Bayesian neural network (BNN) has been widely and successfully employed to describe various nuclear quantities from finite nuclei to infinite nuclear matter, which includes the nuclear mass PhysRevC.93.014311 ; PhysRevC.96.044308 ; NIU201848 ; PhysRevC.100.054311 ; Kejzlar_2020 ; PhysRevC.101.051301 ; PhysRevC.104.014315 ; PhysRevC.106.014305 ; PhysRevC.106.L021303 ; ZHANG2024122820 , $\alpha$-decay properties Rodríguez_2019 ; PhysRevC.107.014310 ; PhysRevC.108.014326 ; PhysRevC.110.024319 and $\alpha$-clustering structure PhysRevC.104.044902 , $\beta$-decay properties PhysRevC.99.064307 ; PhysRevC.106.024306 ; PhysRevC.107.034316 or double-$\beta$-decay models PhysRevD.96.053001 ; PhysRevD.104.055040 ; PhysRevC.94.024603 , $\gamma$-ray emission PhysRevC.105.045801 , nuclear fission Schunck_2020 , low-lying excited phenomena PhysRevC.104.034317 ; WANG2022137154 , landscape of nuclide chart PhysRevC.101.044307 ; PhysRevLett.122.062502 , neutron skin thickness PhysRevC.107.045802 ; PhysRevC.108.024317 , the yields of production cross sections PhysRevC.107.064909 ; PhysRevC.108.044606 ; YingHuaDang , and nuclear charge radii Utama_2016 ; PhysRevC.101.014304 ; PhysRevC.102.054323 ; PhysRevC.105.014308 ; DONG2023137726 , etc. Besides, the Bayesian inference has a guidance to gain deeper insights into the understanding of fundamental interactions PhysRevC.105.044305 ; Wesolowski_2016 ; PhysRevC.97.025805 ; YANG2020135540 ; PhysRevC.103.025807 ; XU2020135820 ; PhysRevC.103.065804 ; PhysRevC.105.014004 , significantly the bulk properties of nuclear matter PhysRevC.96.065805 ; PhysRevLett.125.202702 ; PhysRevC.103.064323 ; PhysRevD.103.063036 ; Xie_2021 ; PhysRevC.105.015806 ; PhysRevC.106.044305 ; PhysRevD.106.043024 ; panglonggang ; PhysRevC.107.045803 ; PhysRevC.108.034908 ; QIU2024138435 ; RADUTA2024138696 ; PhysRevC.107.055803 ; PhysRevC.110.035805 ; PhysRevD.109.123038 . It should be mentioned that Bayesian neural networks can cover the quantitative uncertainties associated with these various physical systems Higdon2015 ; PhysRevC.96.024003 ; PhysRevC.98.024305 ; PhysRevLett.119.252501 ; HU2019134982 ; PhysRevC.101.055803 ; PhysRevC.102.024616 ; PhysRevC.102.054315 ; Lovell_2020 ; PhysRevC.104.054324 ; PhysRevC.106.024607 ; PhysRevC.106.054001 ; PhysRevC.108.024321 ; PhysRevC.109.054301 ; XinyuWang2024 . This encourages us to perform the proceeding researches in exploring the systematical evolution of nuclear charge radii through Bayesian neural network based approach.
@@
Accurate description of nuclear charge radii can be applied to refine the local variations in nuclear size. In Ref. Utama_2016 , it demonstrates that the BNN can increase the prediction ability of nuclear charge radii by training the residuals between the theoretical predictions and experimental data. This leads to the root-mean-square (rms) deviation with being around $0.02$ fm. The improved BNN which considers the pairing and shell closure effects gives the rms deviation $0.015$ fm in calibrating the charge radii data of nuclei with mass number $A\geq 40$ and proton number $Z\geq 20$ PhysRevC.105.014308 . As is well known, the larger odd-even staggering in charge radii can be profoundly observed in the neutron-deficient isotopes around the proton number $Z=82$ isotopic chains Marsh:2018wxs ; PhysRevLett.127.192501 ; PhysRevC.95.044324 ; ANSELMENT1986471 ; PhysRevC.104.024328 ; PhysRevLett.126.032502 . This is attributed to the underlying mechanism of the abrupt shape-phase transition. By considering the abnormal shape staggering effect, the standard rms deviation falls into $0.014$ fm for both the trained and validated charge radii data through the refined Bayesian neural network DONG2023137726 . In addition, different kinds of machine learning methods, such as the naive Bayesian probability classifier PhysRevC.101.014304 and artificial neural networks PhysRevC.102.054323 ; Akkoyun_2013 ; PhysRevC.108.034315 , can be used to explore the systematical evolution of nuclear charge radii as well.
@@
As mentioned above, the Bayesian neural networks can capture the considered error quantification through comparing the experimental data and theoretical simulations. Moreover, this probabilistic model is more computational saving with respect to the generally used energy density functionals (EDFs). As mentioned in Ref. Zhang10662945 , the aleatoric and epistemic uncertainties should be explicitly considered in a given Bayesian neural network. This means that more complex input structures should be taken into account in the simulating process. In other words, more underlying physical mechanisms should be captured as much as possible in the calibrated protocol. To further reduce the quantified uncertainty in validating the nuclear charge radii, both of the quadrupole deformation $\beta_{20}$ and higher moment deformation $\beta_{40}$ are incorporated into the training process in this work. The Casten factor $P$, which is characterized as the correlation between the valence neutrons $\nu_{n}$ and the valence protons $\nu_{p}$, is associated to the shell closure effect PhysRevLett.58.658 ; Angeli_1991 ; Casten_1996 . In our calculations, the Casten factor $P$, valence neutrons $\nu_{n}$, and valence protons $\nu_{p}$ are simultaneously considered as the input parameter sets.
@@
Actually, Casten factor $P$ plays an indispensable role in the calibrated protocol PhysRevC.105.014308 ; DONG2023137726 ; Sheng:2015poa . For the calculation of the Casten factor $P$, the reference neutron and proton magic numbers are taken as $Z=2$, $6$, $14$, $28$, $50$, $82$, $(114)$ and $N=2$, $8$, $14$, $28$, $50$, $82$, $126$, $(184)$ PhysRevC.88.011301 ; Angeli_1991 ; Sheng:2015poa . However, the traditional Casten factor, which is defined as $P=\nu_{n}\nu_{p}/(\nu_{n}+\nu_{p})$ is invalid along $Z=28$, $50$, $82$ isotopic chains due to the fully filled proton magic shells. In order to avoid the aforementioned problem, the shell closure effect has been taken into account through the modified Casten factor $P^{*}$. In this work, the Monte Carlo dropout Bayesian neural network (MC-dropout BNN) approach is built appropriately to describe the charge radii of nuclei with proton number $Z\geq 20$ and mass number $A\geq 40$. The input structures include the proton number ($Z$), mass number ($A$), the pairing effect ($\delta$), isospin effect ($I^{2}$), shell closure effect associated to the Casten factor $P^{*}$, valence neutrons $\nu_{n}$, valence protons $\nu_{p}$, quadrupole deformation $\beta_{20}$, high order hexadecapole deformation $\beta_{40}$, and the local “abnormal” shape staggering effect of 181,183,185Hg ($LI$). The results obtained by incorporating the traditional Casten factor $P$ into the input structures are also shown for comparison.
@@ II Theoretical formalism
The Bayesian neural network (BNN), which is characterized as the combination of an artificial neural network (ANN) and the Bayesian statistical theory, provides a potential approach to comprehend the physical quantities with the associated uncertainty. However, as demonstrated in Ref. JMLRv15srivastava14a , the over-fitting puzzle can be encountered in the traditional neural network. Then the proposed dropout approach can be used to avoid over-fitting problem of the neural network parameters through the regularization technique. Meanwhile, the constructed Monte Carlo dropout (MC-dropout) method is being equivalent to variational inference Bayesian learning method in evaluating a neural network Gal2016 . Therefore the Monte Carlo dropout Bayesian neural network (MC-dropout BNN) model has been built based on these motivations. This approach has been successfully used to analyze various physical quantities, such as the spectral function PhysRevResearch.4.043082 , mass distributions of the induced fission Huo2023 , and electron-carbon scattering data PhysRevC.110.025501 .
@@
where $\sigma$ depicts the nonlinear activation function, $x$ represents the input data, $W_{1}$ represents the weight matrix connecting the input layer to the hidden layer, $b$ denotes the bias vectors, and $W_{2}$ represents the weight matrix that links the hidden to output layer. $z_{1}$ and $z_{2}$ are binary vectors in the sampling process. MC-dropout approach is driven by the Bernoulli distributed random variables in the multiplying hidden activations. The random variables take the value of $1$ with probability parameter $p$ and 00 for other cases Gal2016 . This means the neuron is invalid if the corresponding binary variable takes 00 in a given input.
@@
The first term denotes the Euclidean loss function used in the training and optimization process of neural network. Where $\hat{y}_{n}$ represents the output of a neural network with $L$ layers and $y_{n}$ denotes the target value. $N$ is the number of training data points. A regularization term is often added into the optimization process as shown in the second term. Here, the weight matrixes of dimensions $K_{i}\times{K_{i-1}}$ are represented by $W_{i}$ and the bias vectors of dimensions $K_{i}$ for each layer $i=1,2,\cdots,L$ are denoted by $b_{i}$. The quantity $\lambda_{\mathrm{decay}}$ represents the regularization parameter Huo2023 .
Figure 1: (Color online)Structure of the Monte Carlo dropout Bayesian neural network used in this work. The number of neurons in the input layer is $10$. The $3$ hidden layers are employed, and the number of neurons in the corresponding hidden layer $1$ and $3$ are $28$ and $71$, respectively. The hidden layer $2$, namely the dropout layer, has $300$ neurons. This diagrammatic sketch just shows the dropout hidden layer, in which the different markers denote the random variables take value of $1$ with probability parameter $p$ and 00 for other cases. The number of neurons in the output layer is $1$.
Shown in Refs. Zhang10662945 ; Hüllermeier2021 , the predicted errors of the observables are mainly originated from epistemic uncertainty and aleatoric uncertainty in using machine learning approaches. This means the chosen input structure parameter sets should capture the underlying physical mechanisms as much as possible. In this work, the experimental charge radii data of nuclei with $Z\geq 20$ and $A\geq 40$ are used to train the MC-dropout BNN PhysRevLett.126.032502 ; ANGELI201369 ; LI2021101440 . As mentioned in $Introduction$, the input structures for the MC-dropout BNN are consisted of the proton number ($Z$), mass number ($A$), isospin asymmetry degree ($I^{2}$), pairing effect ($\delta$), Casten factor ($P$), valence neutrons ($\nu_{n}$) and protons ($\nu_{p}$), quadrupole deformation ($\beta_{20}$), the hexadecapole deformation ($\beta_{40}$), and the profound shape-phase transition in 181,183,185Hg ($LI$). The ground state properties about quadrupole and hexadecapole deformation parameters are taken from Ref. MOLLER20161 .
@@
As mentioned before, the traditional Casten factor $P$ is invalid along the proton magic numbers $Z=28$, $50$, $82$ isotopic chains. To avoid this puzzle, the modified Casten factor $P^{*}$ is proposed as follows,
@@
The input parameter sets which includes the traditional Casten factor $P$ cannot describe the shell quenching phenomena around the neutron magic numbers. In contrast to the traditional Casten factor $P$, the modified Casten factor $P^{*}$ can give more high accuracy in the training and validation sets, and the shell closure effect in charge radii can be reproduced well. More details can be found in the following $Section$. For the modified Casten factor $P^{*}$, the input structures can be redefined as $z=\{Z,$ $A$, $I^{2}$, $\delta$, $P^{*}$, $\nu_{n}$, $\nu_{p}$, $\beta_{20}$, $\beta_{40}$, $LI\}$. Meanwhile, if the proton numbers and neutron numbers simultaneously occupy the fully filled shells, such as $(Z=28,N=28)$, $(Z=28,N=50)$, $(Z=50,N=50)$, $(Z=50,N=82)$, and $(Z=82,N=126)$ cases, two kinds of Casten factors are vanished totally. In addition, the np formula Nerlo-Pomorska:1994dhg can also be chosen as our theoretical model to be refined by MC-dropout BNN:

### v2

@@ Abstract
Charge radii can be generally used to encode information about various fine structures of finite nuclei. In this work, a constructed Bayesian neural network based on the Monte Carlo dropout approach is proposed to accurately describe the charge radii of nuclei with proton number $Z\geq 20$ and mass number $A\geq 40$. More motivated underlying mechanisms are incorporated into this combined model in addition to the basic building blocks with the specific number of protons and neutrons, which naturally contain the pairing effect, the isospin effect, the shell closure effect associated with the Casten factor $P$, the valence neutrons, the valence protons, the quadrupole deformation $\beta_{20}$, the high order hexadecapole deformation $\beta_{40}$, and the local shape staggering effect of 181,183,185Hg. To avoid the distorted cases of the traditional Casten factor at the fully filled shells, the modified Casten factor $P^{*}$ is introduced into the input structure parameter sets. The standard root-mean-square deviation is reduced to $0.0084$ fm for the training data set and $0.0124$ fm for the validation data set with the modified Casten factor $P^{*}$. Meanwhile, the shell closure effect of nuclear charge radii can be reproduced remarkably well. We have successfully demonstrated the ability of this constructed model to significantly increase the accuracy in predicting the nuclear charge radii.
@@ I Introduction
Finite nuclei, characterized as strongly correlated many-body systems, provide a natural probe for investigating fundamental interactions and forces. As directly measured quantities in atomic nuclei, charge radii can be generally used to reflect the range of the proton density distributions and encode information about various fine structure phenomena, such as the emergence of shell closure effects Angeli_2015 ; PhysRevC.100.044310 ; GarciaRuiz:2019cog ; PhysRevC.104.064313 ; PhysRevLett.129.142502 and the odd-even staggering (OES) behavior GarciaRuiz:2016ohj ; Marsh:2018wxs ; Miller2019 ; PhysRevLett.127.192501 . Furthermore, the available description of the proton density distributions can also be used to pin down the components of the isospin interactions in the equation of state (EoS) of asymmetric nuclear matter PhysRevC.88.011301 ; PhysRevLett.119.122502 ; PhysRevResearch.2.022035 ; PhysRevLett.127.182503 ; nuclscitech34.119 ; PhysRevLett.132.162502 . This means that more data on nuclear charge radii should be urgently required with highly accurate precision. Remarkable achievements in detecting charge radii have been achieved over the past decades. Especially for nuclei with extreme neutron-to-proton ratios, the laser spectroscopy method provides a credible access to measure the nuclear size CAMPBELL2016127 ; YANG2023104005 .
@@
The undertaken efforts have been devoted to featuring the normal and abnormal variations in nuclear charge radii along a long isotopic family, which includes the empirical formulas Zhang2002EPJA285 ; PhysRevC.87.024310 ; Piekarewicz2010 ; PhysRevC.94.064315 , relativistic mean field theory Geng:2003pk ; PhysRevC.102.024307 ; PhysRevC.109.064302 , non-relativistic Skyrme Hartree-Fock-Bogoliubov (HFB) approach PhysRevLett.102.242501 ; PhysRevC.82.035804 , and the Fayans energy density functional PhysRevC.95.064328 ; PhysRevLett.122.192502 , etc. Recently, machine learning methods have been made great applications in the course of nuclear physics PhysRevC.98.034318 ; Bedaque2021 ; Paquet_2024 . Particularly, a well-known approach based on Bayesian neural network (BNN) has been widely and successfully employed to describe various nuclear quantities from finite nuclei to infinite nuclear matter, which includes the nuclear mass PhysRevC.93.014311 ; PhysRevC.96.044308 ; NIU201848 ; PhysRevC.100.054311 ; Kejzlar_2020 ; PhysRevC.101.051301 ; PhysRevC.104.014315 ; PhysRevC.106.014305 ; PhysRevC.106.L021303 ; ZHANG2024122820 , $\alpha$-decay properties Rodríguez_2019 ; PhysRevC.107.014310 ; PhysRevC.108.014326 ; PhysRevC.110.024319 and $\alpha$-clustering structure PhysRevC.104.044902 , $\beta$-decay properties PhysRevC.99.064307 ; PhysRevC.106.024306 ; PhysRevC.107.034316 or double-$\beta$-decay models PhysRevD.96.053001 ; PhysRevD.104.055040 ; PhysRevC.94.024603 , $\gamma$-ray emission PhysRevC.105.045801 , nuclear fission Schunck_2020 , low-lying excited phenomena PhysRevC.104.034317 ; WANG2022137154 , landscape of nuclide chart PhysRevC.101.044307 ; PhysRevLett.122.062502 , neutron skin thickness PhysRevC.107.045802 ; PhysRevC.108.024317 , the yields of production cross sections PhysRevC.107.064909 ; PhysRevC.108.044606 ; YingHuaDang , and nuclear charge radii Utama_2016 ; PhysRevC.101.014304 ; PhysRevC.102.054323 ; PhysRevC.105.014308 ; DONG2023137726 , etc. In addition, Bayesian inference can help to gain deeper insights into the understanding of fundamental interactions PhysRevC.105.044305 ; Wesolowski_2016 ; PhysRevC.97.025805 ; YANG2020135540 ; PhysRevC.103.025807 ; XU2020135820 ; PhysRevC.103.065804 ; PhysRevC.105.014004 , significantly the bulk properties of nuclear matter PhysRevC.96.065805 ; PhysRevLett.125.202702 ; PhysRevC.103.064323 ; PhysRevD.103.063036 ; Xie_2021 ; PhysRevC.105.015806 ; PhysRevC.106.044305 ; PhysRevD.106.043024 ; panglonggang ; PhysRevC.107.045803 ; PhysRevC.108.034908 ; QIU2024138435 ; RADUTA2024138696 ; PhysRevC.107.055803 ; PhysRevC.110.035805 ; PhysRevD.109.123038 . It should be mentioned that Bayesian neural networks can cover the quantitative uncertainties associated with these various physical systems Higdon2015 ; PhysRevC.96.024003 ; PhysRevC.98.024305 ; PhysRevLett.119.252501 ; HU2019134982 ; PhysRevC.101.055803 ; PhysRevC.102.024616 ; PhysRevC.102.054315 ; Lovell_2020 ; PhysRevC.104.054324 ; PhysRevC.106.024607 ; PhysRevC.106.054001 ; PhysRevC.108.024321 ; PhysRevC.109.054301 ; XinyuWang2024 . This encourages us to perform the proceeding research in exploring the systematic evolution of nuclear charge radii through Bayesian neural network-based approach.
@@
An accurate description of nuclear charge radii can be applied to refine the local variations in nuclear size. As demonstrated in Ref. Utama_2016 , BNN enhances the prediction ability of nuclear charge radii by training the residuals between the theoretical predictions and the experimental data. This leads to the deviation of the root-mean-square (rms) around $0.02$ fm. The improved BNN that considers the pairing and shell closure effects gives the rms deviation $0.015$ fm in calibrating the charge radii data of nuclei with mass number $A\geq 40$ and proton number $Z\geq 20$ PhysRevC.105.014308 . As is well known, the larger odd-even staggering in the charge radii can be profoundly observed in the neutron-deficient isotopes around the proton number $Z=82$ isotopic chains Marsh:2018wxs ; PhysRevLett.127.192501 ; PhysRevC.95.044324 ; ANSELMENT1986471 ; PhysRevC.104.024328 ; PhysRevLett.126.032502 . This is attributed to the underlying mechanism of the abrupt shape-phase transition. Considering the abnormal shape-staggering effect, the standard rms deviation falls into $0.014$ fm for both the trained and validated charge radii data through the refined Bayesian neural network DONG2023137726 . In addition, different kinds of machine learning methods, such as the naive Bayesian probability classifier PhysRevC.101.014304 and artificial neural networks PhysRevC.102.054323 ; Akkoyun_2013 ; PhysRevC.108.034315 , can be used to explore the systematic evolution of nuclear charge radii as well.
@@
As mentioned above, Bayesian neural networks can capture the considered error quantification by comparing the experimental data and theoretical simulations. As mentioned in Ref. Zhang10662945 , the aleatoric and epistemic uncertainties should be explicitly considered in a given Bayesian neural network. This means that more underlying physical mechanisms should be captured as much as possible in the calibrated protocol. To further reduce the quantified uncertainty in validating the nuclear charge radii, both quadrupole deformation $\beta_{20}$ and higher moment deformation $\beta_{40}$ are incorporated into the simulated codes in this work. The Casten factor $P$, which is characterized as the correlation between the valence neutrons $\nu_{n}$ and the valence protons $\nu_{p}$, is associated with the shell closure effect PhysRevLett.58.658 ; Angeli_1991 ; Casten_1996 . In our calculations, the Casten factor $P$, valence neutrons $\nu_{n}$, and valence protons $\nu_{p}$ are simultaneously considered as input parameter sets.
@@
Actually, the Casten factor $P$ plays an indispensable role in the calibrated protocol PhysRevC.105.014308 ; DONG2023137726 ; Sheng:2015poa . For the calculation of the Casten factor $P$, the reference neutron and proton magic numbers are taken as $Z=2$, $6$, $14$, $28$, $50$, $82$, $(114)$ and $N=2$, $8$, $14$, $28$, $50$, $82$, $126$, $(184)$ PhysRevC.88.011301 ; Angeli_1991 ; Sheng:2015poa . However, the traditional Casten factor, which is defined as $P=\nu_{n}\nu_{p}/(\nu_{n}+\nu_{p})$, is distorted due to the fully filled proton magic shells or the neutron magic numbers. In order to avoid the aforementioned problem, the shell closure effect has been taken into account by using the modified Casten factor $P^{*}$. In this work, the Monte Carlo dropout Bayesian neural network approach (MC-dropout BNN) is built appropriately to describe the charge radii of nuclei with proton number $Z\geq 20$ and mass number $A\geq 40$. The input structures include the proton number ($Z$), the mass number ($A$), the pairing effect ($\delta$), the isospin effect ($I^{2}$), the shell closure effect associated with the modified Casten factor $P^{*}$, the valence neutrons $\nu_{n}$, the valence protons $\nu_{p}$, the quadrupole deformation $\beta_{20}$, the high order hexadecapole deformation $\beta_{40}$ and the local ‘abnormal’ shape staggering effect of 181,183,185Hg ($LI$). The results obtained by incorporating the traditional Casten factor $P$ into the input structures are also shown for comparison.
@@ II Theoretical formalism
The Bayesian neural network (BNN), which is characterized as the combination of an artificial neural network (ANN) and Bayesian statistical theory, provides a potential approach to comprehend the physical quantities with the associated uncertainty. However, as demonstrated in Ref. JMLRv15srivastava14a , the over-fitting puzzle can be encountered in the traditional neural network. Then the proposed dropout approach can be used to avoid over-fitting problem of the neural network parameters through the regularization technique. Meanwhile, the constructed Monte Carlo dropout method (MC-dropout) is being equivalent to variational inference Bayesian learning method in evaluating a neural network Gal2016 . Moreover, this probabilistic model is more time saving compared to the generally used energy density functionals (EDFs). In contrast to the traditional BNN, the Monte Carlo dropout Bayesian neural network (MC-dropout BNN) model can save more computing resources as well. Therefore, the MC-dropout BNN model has been built based on these motivations. This approach has been successfully used to analyze various physical quantities, such as the spectral function PhysRevResearch.4.043082 , mass distributions of the induced fission Huo2023 , and the electron-carbon scattering data PhysRevC.110.025501 .
@@
where $\sigma$ depicts the nonlinear activation function, $x$ represents the input data, $W_{1}$ represents the weight matrix connecting the input layer to the hidden layer, $b$ denotes the bias vectors, and $W_{2}$ represents the weight matrix that links the hidden to output layer. $z_{1}$ and $z_{2}$ are binary vectors in the sampling process. MC-dropout approach is driven by the Bernoulli distributed random variables in the multiplying hidden activations. The random variables take the value of $1$ with probability parameter $p$ and 00 for other cases gal16 ; Gal2016 . This means that the neuron is invalid if the corresponding binary variable takes 00 in a given input. In our calculations, the dropout probability parameter is chosen as 0.5.
@@
The first term denotes the Euclidean loss function used in the neural network training and optimization process. Where $\hat{y}_{n}$ represents the output of a neural network with $L$ layers and $y_{n}$ denotes the target value. $N$ is the number of training data sets. A regularization term, as shown in the second term, is often added to the optimization process. Here, the weight matrices of the dimensions $K_{i}\times{K_{i-1}}$ are represented by $W_{i}$ and the bias vectors of the dimensions $K_{i}$ for each layer $i=1,2,\cdots,L$ are indicated by $b_{i}$. The hyperparameter $\lambda_{\mathrm{decay}}$ in Eq. (2) represents the regularization parameter Huo2023 , which is used to reduce the risk of overfitting and its value is $3\times 10^{-9}$. Meanwhile, the hyperparameters used in this work are identical for the constructed models (Fig. 1).
Figure 1: Structure of the Monte Carlo dropout Bayesian neural network used in this work. The number of neurons in the input layer is $8$ or $10$. The $3$ hidden layers are employed, and the number of neurons in the corresponding hidden layer $1$ and $3$ are $28$ and $71$, respectively. The hidden layer $2$, namely the dropout layer, has $300$ neurons. This diagrammatic sketch just shows the dropout hidden layer, in which the different markers denote the random variables take value of $1$ with probability parameter $p$ and 00 for other cases. The number of neurons in the output layer is $1$.
Shown in Refs. Zhang10662945 ; Hüllermeier2021 , the predicted errors of the observables are mainly originating from epistemic uncertainty and aleatoric uncertainty in using machine learning approaches. This means that the chosen input structure parameter sets should capture the underlying physical mechanisms as much as possible. In this work, the experimental charge radii data of nuclei with $Z\geq 20$ and $A\geq 40$ are used to train the MC-dropout BNN PhysRevLett.126.032502 ; ANGELI201369 ; LI2021101440 . As mentioned in $Introduction$, the input structures for the MC-dropout BNN are consisted of the proton number ($Z$), mass number ($A$), isospin asymmetry degree ($I^{2}$), pairing effect ($\delta$), Casten factor ($P$), valence neutrons ($\nu_{n}$) and protons ($\nu_{p}$), quadrupole deformation ($\beta_{20}$), the hexadecapole deformation ($\beta_{40}$), and the profound shape-phase transition in 181,183,185Hg ($LI$). The ground-state properties about the quadrupole and hexadecapole deformation parameters are taken from Ref. MOLLER20161 .
@@
As mentioned above, the traditional Casten factor $P$ is invalid along the isotopic chains of the proton magic numbers $Z=28$, $50$, and $82$ or the isotonic chains with the neutron numbers $N=28$, $50$, $82$, and $126$. To avoid this puzzle, the modified Casten factor $P^{*}$ is proposed as follows,
@@
The input parameter sets that include the traditional Casten factor $P$ cannot describe the shell quenching phenomena well. In contrast to the traditional Casten factor $P$, the modified Casten factor $P^{*}$ can give more high accuracy in the training and validation sets, and the shell closure effect in the charge radii can be reproduced well. More details can be found in the following $Section$. For the modified Casten factor $P^{*}$, the input structures can be redefined as $z=\{Z,$ $A$, $I^{2}$, $\delta$, $P^{*}$, $\nu_{n}$, $\nu_{p}$, $\beta_{20}$, $\beta_{40}$, $LI\}$. Meanwhile, if the proton numbers and neutron numbers simultaneously occupy the fully filled shells, such as cases $(Z=28,N=28)$, $(Z=28,N=50)$, $(Z=50,N=50)$, $(Z=50,N=82)$, and $(Z=82,N=126)$, two kinds of Casten factor vanish completely. In addition, the np formula Nerlo-Pomorska:1994dhg can also be chosen as our theoretical model to be refined by MC-dropout BNN: